OpenStudy (anonymous):

A trapezoid has an area of 22 square inches. The longer base measures 6 inches and the height is 4 inches. use the formula

OpenStudy (anonymous):

6 times 4=24

OpenStudy (anonymous):

A = [(B1 + B2) * H]/2 22 = [(B1 + 6) * 4]/2 44 = 4B1 + 24 20 = 4B1 5 = B1